# Towards Identifiability of Interventional Stochastic Differential Equations

### v2

@@ Abstract
We study identifiability of stochastic differential equation (SDE) models under multiple interventions. Our results give the first provable bounds for unique recovery of SDE parameters given samples from their stationary distributions. We give tight bounds on the number of necessary interventions for linear SDEs, and upper bounds for nonlinear SDEs in the small noise regime. We experimentally validate the recovery of true parameters in synthetic data, and motivated by our theoretical results, demonstrate the advantage of parameterizations with learnable activation functions.
@@ 1 Introduction
High-throughput perturbations, e.g., using CRISPR systems, coupled with single-cell genomic sequencing provide the opportunity to study the dynamic and causal relationships (known as the “gene regulatory network”, GRN) between genes at an unprecedented resolution, under a wide set of possible interventions, informing which genes should be targeted for treatment [12]. While substantial progress has been made in applying machine learning and causal inference methods to such genomic data, there has been limited theoretical work to understand when such models are identifiable.
@@
Due to the high degree of gene co-expression, substantial noise and unmeasured modalities (e.g., chromatin state, protein expression), latent confounding is particularly severe in genomic data, leading to considerable false positive rates among popular methods [20]. Causal disentanglement aims to learn causal factors in spite of these confounders, mainly focusing on directed acyclic graph (DAG) based methods. To demonstrate these methods are well-founded, there is considerable effort devoted to understanding which models have identifiability guarantees [23]. However, these models suffer from inherent weakness, in particular 1) being unable to represent cycles (a common occurrence in GRNs) or 2) approximate continuous-time dynamical models.
@@
There has been renewed interest in inferring GRNs directly with dynamical systems [31]. There is precedent for this type of modeling to represent the so- called “Waddington landscape” [41], the hypothetical energy surface that cells traverse as they differentiate. Furthermore, SDEs are commonly used for simulating transciptomic datasets from a given gene regulatory network [32, 11], to benchmark GRN recovery algorithms.
As recently expounded in the context of diffusion models (which can be thought of as time-discretized SDEs), SDEs can fit observational data [39]. But without a restriction on the model, it is impossible to identify the true SDE that will generalize to new interventions, either on unseen gene perturbations or intervening on several genes simultaneously. Furthermore, dynamical systems typically learn from many trajectories. But the destructive nature of sequencing makes it impossible to observe the trajectory of a single cell at multiple time-points, and in general it is difficult to obtain any time-series genomics data due to the high expense. Therefore, practitioners often only collect data at the end of an experiment, i.e., from the stationary distribution of the system. Consequently, several dynamics-based models rely on single time points for evaluation, and this lack of trajectories may render the true biological model unidentifiable.
@@
Our interest in this work is to verify which parametric assumptions are necessary for dynamical systems to have these guarantees. Namely:
@@
In this work, we offer the first analysis of identifiability of stochastic differential equations with interventions. Specifically:
@@
  • We apply this insight to synthetic and semi-synthetic data to confirm the efficacy of learned activations in causal SDEs, which improve expressiveness without sacrificing a simple structure from which to extract a gene regulatory network.
@@ 2.1 Notation
[…]
@@ 2.2 Stochastic Differential Equations
We will only consider autonomous systems, i.e. where the drift and noise terms have no dependence on time $t$. We will enforce the weak conditions on the drift and noise to guarantee a unique stationary distribution [4], with a density $p$ that satisfies the Fokker-Planck equation,
@@ 2.4 Interventional SDEs
We focus on the setting where we only observe the SDE through the induced stationary density under $\mathcal{I}$ different shift interventions. Specifically, there are vectors $\{c_{i}\}_{i=1}^{\mathcal{I}}$ with each $c_{i}\in\mathbb{R}^{n}$, and we observe the stationary distribution of the SDE,
@@
In the gene regulation setting, $c$ could represent CRISPR activation (one element of $c$ being positive) or interference (one element being negative) to enhance or repress gene expression, respectively.
We denote the concatenated intervention column vectors by the matrix $C\in\mathbb{R}^{n\times\mathcal{I}}$. These shifts could correspond to overexpression or knockdown in Perturb-seq data [12]. One appeal of dynamical systems over DAGs is that soft interventions such as these may be expressed very naturally with a shift vector added to the drift in the SDE, while a DAG cannot capture how a feedback loop might be changed through overexpression or knockdown.
In the causal disentanglement literature, soft interventions typically give fewer guarantees [6, 40]. Therefore, we assume knowledge of the interventions,
@@ 3.1 Causal Representation
Genomics is only one motivation for a formal model of causality [29], a field far too broad to thoroughly survey here. The most popular underlying model is the structural causal model (SCM), which characterizes the conditional distribution of a random variable under arbitrary intervention. Learning an SCM typically requires very strong assumptions such as sparsity [38], interventional data [23], parametric assumptions [30], among many other results. Sparsity is a very common theme in these models, though it may also be expressed in an assumption that the number of latent variables is small, i.e. a low-rank constraint [13].
@@ 3.2 Causal Disentanglement with Interventional Data
The bulk of the literature on causal disentanglement focuses on SCMs with an underlying DAG. Particularly relevant to our work are results that assume access to multiple interventional environments, either acting directly on the observed variables [5] or identifying a latent model under some distributional assumptions [23, 40, 6].
Some of these works have addressed the crucial limitation of acyclicity [45, 24, 2], but without necessarily incorporating dynamics. Many works require hard interventions, where an intervened variable is a function of exogenous noise, although some can handle soft interventions [44].
@@ 3.3 Dynamical System Methods
Several authors have considered modeling perturbations with dynamical systems [31]. In particular, these models assume observed data is drawn from the stationary distribution under an SDE, and match a learned SDE under numerous interventional regimes. These works focus on linear drift and intervention-dependent parameters [33] and non-linear drift with shift interventions [26]. Notably, neither paper gives theory to confirm whether these interventional models are identifiable.
Nevertheless, works drawing on dynamical systems are regularly applied to the task of inferring GRNs. Some methods act on pseudotime, an inferred notion of time from cell states that enables the machinery of dynamical systems even when very few “real” timepoints are available [42, 18]. Although harder to obtain due the destructive nature of RNA sequencing, genuine temporal data (even with very few timepoints) can also be modeled with the intent of extracting a GRN [25].
The closest work to ours proves an identifiability bound for linear SDEs under a strong sparsity assumption [10]. However, their result doesn’t consider any interventional data, and the exact pattern of sparsity in the drift matrix is assumed to be known a priori, which is rarely the case in applied problems of interest. Another closely related work is [17], which can simultaneously infer the drift and diffusion. However, this work differs from our setting in that it assumes linearity, doesn’t study interventions, and works with multiple temporal marginals rather than just the stationary distribution.
@@ 4 Main Results
We consider two main parameterizations, whether the drift is linear or a two-layer neural network (an MLP), and in which regimes the parameters may be uniquely identified up to appropriate invariances. Prior work has considered sparsity in the linear setting only [10], but we focus on parameterizations that project to a low-dimensional space, through small rank or hidden dimension. In other words, although the ambient dimension $n$ may be large, we assume the hidden dimension $r$ is much smaller, and ideally the number of interventions to uniquely identify the parameters should scale with $r$. In the genomics context, $n$ is the number of modeled (e.g., “highly variable”) expressed genes in the thousands, and $r$ corresponds to a much smaller number of co-expressed gene modules/pathways.
@@ 4.1 Linear case
We start with the linear case and consider the parameterization $v(x)=(AB-D)x$ where $A\in\mathbb{R}^{n\times r}$, $B\in\mathbb{R}^{r\times n}$, $D\in\mathbb{R}^{n\times n}$. Furthermore, to ensure the matrix is Hurwitz and the SDE therefore has a stationary distribution, we assume $\|AB\|\leq\gamma<1$ and $D\succeq I$. We associate $AB$ with transcriptional regulation, and $D$ with decay of each RNA species.
This parameterization is a simplification of the one used in [33], which also uses linear SDEs but considers hard interventions with intervened genes having entirely new rows in the drift matrix, rather than shift interventions. In our biological context, the low rank assumption is 1) intuitive in that genes are expected to behave in coordinated ways in modules/pathways, and 2) empirically proven, in that almost all successful single-cell methods leverage dimensionality reduction.
@@
In the linear setting, there exist choices for parameters $A,B,D$ and interventions $C$ such that the parameters are not identifiable with less than $n-r$ interventions.
The proof is provided in Appendix A. The intuition is that one can choose interventions that don’t affect the system dynamics. But this situation is pathological, and in the linear case under some weak assumptions about the distribution of the parameters and the interventions we have identifiability. We start with assumptions about the distribution of the SDE parameters and the interventions:
@@
The rotational invariance assumption on the distributions of $A$ and $B$ encodes an uninformative prior on which low-rank subspace governs the dynamics. Indeed, a simple choice that satisfies this assumption is sampling $A$ and $B^{T}$ from the uniform measure on the Stiefel manifold (the space of rectangular matrices with orthonormal columns) and then scaling down slightly to ensure a spectral norm strictly smaller than one.
Additionally, for theoretical tractability, we presume knowledge of the decay term $D$. This is a stronger assumption, but defendable in our genomics context. Typically, we’re more interested in dynamics of genes regulating each other rather than self-regulation, and furthermore from a systems perspective there is no way to distinguish self-regulation from RNA decay. While decay rate varies across genes, it is less regulated by other genes than the transcription rate, so external experimental estimates of a diagonal $D$ (e.g., from BRIC-seq [19]) could be used. Hence, fixing the decay rate leaves the focus of learning on the presumed low-rank dynamics.
@@
The true decay matrix $D$ is observed.
Altogether, we can now present a tight identifiability result in the linear case.
@@
Consider the linear drift in Equation 6. Suppose $A$, $B$ are drawn from any density further constrained such that $\|A\|,\|B\|\leq\gamma<1$. Then the drift $AB-D$ is identifiable a.s. with $r$ interventions, and unidentifiable a.s. with at most $r-2$ interventions.
@@
If diagonal decay term $D$ is not inferred in advance, learning it simultaneously is comparable to the setting of robust PCA [8] or recovery of a diagonal plus a positive semidefinite low rank term [37]. However, unlike the usual matrix completion setting where entries of the matrix are revealed uniformly at random, we have access to correlated low-rank measurements of the form $e_{i}c_{j}^{T}$ for $\{e_{i}\}_{i=1}^{n}$ the elementary basis. This setting is well studied with sub-Gaussian measurement vectors [46], but the tools do not readily apply to our setting with non-random measurements and the constraints of the Lyapunov equation. Nevertheless, we observe in Section 5 that identifiability even with an unknown diagonal decay matrix $D$ is empirically achievable in this setting, while still subject to the lower bound established in Theorem 4.4.
@@ 4.2 Nonlinear case
The nonlinear case is more challenging, primarily because there is no longer a (nearly) closed form characterization of the stationary distribution. To apply tools from above, we consider a regime where a linearized SDE can approximately capture the true dynamics, by restricting to contractive drift and small noise.
In choosing a sensible drift parameterization, we want to balance expressiveness against low depth, as in single cell applications the induced modules that drive cell evolution are ideally interpretable in their own right, as well as requiring the induced SDE to have a stationary distribution at all.
@@
The strongest constraint here is condition 1, upper bounding the first derivative of the activation, as it implies the noiseless dynamics are globally contractive, but it guarantees a stationary distribution exists for any intervention vector $c$. In practice, cells do not (in fact, cannot) produce unlimited amounts of RNA, so conditions that ensure the existence of a stationary distribution can be motivated from the biological perspective. Furthermore, these conditions still allow for a wide class of possible activations $\sigma$, mainly constrained to be increasing with bounded first and second derivative. We observe the improved expressiveness of the stationary distribution of nonlinear SDEs in Figure 1. Note constraint 4 rules out functions that are linear (or locally linear). This enables a stronger possible identifiability guarantee than only recovering $A$ and $B$ up to rotation as in the linear case.
Figure 1: Contour plot of the stationary SDE under different activations, namely observed linear (top left), intervened linear (bottom left), observed non-linear (top right), and intervened non-linear (bottom right). While the activation contractivity enforces one mode, the linear distribution is Gaussian with fixed covariance across interventions, while the nonlinear distribution can vary more.
@@
Setting $\epsilon=0$ in equation 6, the parameters are a.s. not identifiable with fewer than $n-r$ interventions.
The proof is in Appendix A. Intuitively, the equilibrium points of the noiseless ODE approximate the mean of the SDE for small noise (note this is not true for larger noise, for example see [28]). This suggests that in order to make use of the low-rank assumption on the dynamics, one must at least take advantage of second order moments of the SDE even in the small noise limit.
@@
Let $x^{*}$ be the unique point satisfying $v(x^{*})=0$, and define $L=Jv(x^{*})$. If $\omega$ is the solution to the Lyapunov equation $L\omega+\omega L^{T}+I=0$, and $m_{\epsilon}$ and $\Sigma_{\epsilon}$ are the mean and covariance of the stationary distribution of Equation (6), we have for sufficiently small $\epsilon$,
@@
$\|\Sigma_{\epsilon}/\epsilon-\omega\|$ $\lesssim\frac{\epsilon^{1/2}r^{2}n^{1/2}\sqrt{M}}{(1-\gamma)^{3}}.$ (8)
The proof is given in Appendix A. This is one instance of perturbation theory for SDEs [15, 35].
Equipped with this fact, one can consider access to the stationary distribution of an SDE, specifically the first and second-order moments, and inspect what happens now as the noise goes to zero. In this case, we obtain non-trivial identifiability, i.e. a bound that scales with the hidden dimension $r$:
@@
We give a brief sketch of the proof, to capture the novel elements and how the identifiability appears. The zero-noise limit allows us to linearize the SDE under different interventions. However, unlike the truly linear case, the covariance is not identical among all interventions. By Theorem 4.7, the covariance of the $i$th intervention approximately corresponds to the SDE with drift $A(J\sigma(Bx_{i}^{*}))B$ where $x_{i}^{*}$ is the unique zero of the true drift $v(x)+c_{i}$. Sufficient variability in these fixed points enables the identification of $A$ and $B$ by taking linear combinations of these drift matrices. Because the Jacobian term is diagonal, a linear combination of drifts across $r$ interventions can yield a rank-one matrix, which corresponds to simultaneously recovering a column of $A$ and a row of $B$. But we don’t have access to the drift matrix itself, only the covariance it induces, and so the proof relies on finding a different set of matrices where rank-one elements can be identified.
Crucially, one can verify the rank of linear combinations of matrices without knowledge of the activation function itself. This enables the identification of the bias term, but furthermore one can identify the parameters $A$ and $B$ even if each elementwise activation in $\sigma$ is distinct. This observation motivates learnable activations, which we experiment with in Section 5. In our setting where two-layer MLPs with small width offer better interpretability, parameterizing the activation is an effective way to increase expressiveness without sacrificing identifiability.
We hypothesize that the constraint that the SDE drift have a single absorbing point is likely unnecessary. If the ODE given by $\frac{dx}{dt}=v(x)+c_{i}$ had multiple stable equilibria, then in the limit of small noise, the SDE stationary distribution should approach a Gaussian mixture centered around each equilibrium point. The proof only requires certain linear independence conditions of vectors induced by each fixed point, so multiple equilibria and the local covariance around them would offer essentially the same information as additional interventions. This is consistent with the observations in [26] that nonlinear SDE models generalize well with a limited number of interventions.

### v3

@@ Abstract
We study identifiability of stochastic differential equations (SDE) under multiple interventions. Our results give the first provable bounds for unique recovery of SDE parameters given samples from their stationary distributions. We give tight bounds on the number of necessary interventions for linear SDEs, and upper bounds for nonlinear SDEs in the small noise regime. We experimentally validate the recovery of true parameters in synthetic data, and motivated by our theoretical results, demonstrate the advantage of parameterizations with learnable activation functions in application to gene regulatory dynamics.
@@ 1 Introduction
Stochastic dynamical systems are ubiquitous as models for natural data. They are perfectly suited for application to time-series data, and therefore also a good candidate to characterize systems that reach a steady state in the limit. If a system is governed by some stochastic differential equation (SDE) and the same system is observed under different interventions, ideally one would learn the underlying parameters governing the dynamics, and guarantee accurate prediction under new interventions.
However, in many natural settings, data is modeled as following and SDE even if one does not have access to explicit trajectories. Studies of ecological systems focus on the long-term survival of multiple species modeled by the quasi-stationary state of SDEs with environmental factors as perturbations (Hening & Li, 2021). The application of flow cytometry to protein signaling networks under perturbation (Sachs et al., 2005) is destructive and yields protein quantification at one time point, modeled using the stationary distributions of linear SDEs in Varando & Hansen (2020).
One highly motivating application is single-cell genomic sequencing with high-throughput CRISPR perturbations. Biologists are often interested in inferring the gene regulatory network (GRN) that characterizes the dynamics of gene expression, informing which genes should be targeted for treatment (Dixit et al., 2016). But the destructive nature of sequencing makes it impossible to observe the trajectory of a single cell at multiple time-points, and in general it is difficult to obtain any time-series genomics data due to the high expense. Therefore, practitioners often only collect data at the end of an experiment, i.e., from the stationary distribution of the system.
@@
Understanding the dynamics is essential for extrapolating to unseen settings, but noise and latent confounding makes it non-trivial to determine the true dynamics. Causal disentanglement aims to learn causal factors in spite of these confounders, mainly focusing on directed acyclic graph (DAG) based methods. To demonstrate these methods are well-founded, there is considerable effort devoted to understanding which models have identifiability guarantees (Lachapelle et al., 2022). However, these models suffer from inherent weakness, in particular 1) being unable to represent cycles or 2) approximate continuous-time dynamical models.
@@
There has been renewed interest in modeling with stochastic differential equations (SDE) directly (Peters et al., 2022). In the genomic context, there is precedent for this type of modeling to represent the so- called “Waddington landscape” (Waddington, 2014), the hypothetical energy surface of cells. Furthermore, SDEs are commonly used for simulating transcriptomic datasets from a given gene regulatory network (Pratapa et al., 2020; Dibaeinia & Sinha, 2020).
As demonstrated in the context of diffusion models, SDEs are fully expressive in practice and can accurately generate observational data (Song et al., 2021). But to identify the true underlying SDE requires assumptions on the model. Foundational theoretical works on identifying dynamical systems typically learn from many trajectories or even the infinitesimal generator of the dynamics (Hansen & Sokol, 2014), leaving open the harder setting of observing only the stationary distribution.
@@
Our interest in this work is to verify which parametric assumptions are necessary for dynamical systems to have identifiability guarantees without trajectory data. Namely:
@@
In this work, we offer the first analysis of identifiability of interventional stochastic differential equations, with data restricted to the stationary measure. Specifically:
@@
  • We apply this insight to synthetic data and semi-synthetic genomic data to confirm the efficacy of learned activations in causal SDEs, which improve expressiveness without sacrificing a simple structure and enable the inference of gene regulatory networks.
@@ 2.1 Notation
We let $P_{A}$ denote the orthogonal projection onto the image of matrix $A$, and $P_{A}^{\perp}=I-P_{A}$ the orthogonal projection onto its complement. We let $A^{\dagger}$ denote the pseudoinverse of $A$, and note that if $A$ has linearly independent columns it is a left inverse such that $A^{\dagger}A=I$, likewise for rows and right inverse. We let $\|\cdot\|$ denote the spectral norm and $\|\cdot\|_{F}$ the Frobenius norm. We will use $\lesssim$ to denote inequality up to constant factor, and iid as independent and identically distributed.
@@ 2.2 Stochastic Differential Equations
We will only consider autonomous systems, i.e. where the drift and noise terms have no dependence on time $t$. We will enforce the weak conditions on the drift and noise to guarantee a unique stationary distribution (Berglund, 2021), with a density $p$ that satisfies the Fokker-Planck equation,
@@ 2.4 Interventional SDEs
We focus on the setting where we only observe the SDE through the induced stationary density under $k$ different shift interventions. Specifically, there are vectors $\{c_{i}\}_{i=1}^{k}$ with each $c_{i}\in\mathbb{R}^{n}$, and we observe the stationary distribution of the SDE,
@@
We denote the concatenated intervention column vectors by the matrix $C\in\mathbb{R}^{n\times k}$. In the causal disentanglement literature, shift interventions typically give fewer guarantees (Buchholz et al., 2024; Squires et al., 2023). Therefore, we assume knowledge of the interventions:
@@ 3.1 Causal Representation
In terms of modeling causality (Pearl, 2009), the most popular underlying model is the structural causal model (SCM), which characterizes the conditional distribution of a random variable under arbitrary intervention. Learning an SCM typically requires very strong assumptions such as sparsity (Schölkopf et al., 2021), interventional data (Lachapelle et al., 2022), parametric assumptions (Peters & Bühlmann, 2014), among many other results. Sparsity is a very common theme in these models, though it may also be expressed in an assumption that the number of latent variables is small, i.e. a low-rank constraint (Fang et al., 2023).
@@ 3.2 Causal Disentanglement with Interventional Data
The bulk of the literature on causal disentanglement focuses on SCMs with an underlying DAG. Relevant to our work are results that assume access to multiple interventional environments, either acting directly on the observed variables (Brouillard et al., 2020) or identifying a latent model under some distributional assumptions (Lachapelle et al., 2022; Squires et al., 2023; Buchholz et al., 2024).
Some of these works have addressed the crucial limitation of acyclicity (Zheng et al., 2018; Lee et al., 2019; Atanackovic et al., 2023), but without necessarily incorporating dynamics. Many works require hard interventions, where an intervened variable is a function of exogenous noise, although some can handle soft interventions (Zhang et al., 2024).
@@ 3.3 Dynamical System Methods
Previous work has considered modeling perturbations with dynamical systems (Peters et al., 2022). One can prove identifiability of SDE parameters from the generator (Hansen & Sokol, 2014) or trajectories (Guan et al., 2024). Other methods work in our harder setting where observed data is drawn from the stationary distribution under an SDE, and match a learned SDE under numerous interventions. These works focus on linear drift and intervention-dependent parameters (Rohbeck et al., 2024) or non-linear drift with shift interventions (Lorch et al., 2024). Notably, neither paper gives theory to confirm if these models are identifiable.
There are also methods specific to a particular scientific domain. In genomics, some methods act on pseudotime, an inferred notion of time from cell states when very few “real” timepoints are available (Wang et al., 2024; Hossain et al., 2024). Although harder to obtain due the destructive nature of RNA sequencing, genuine temporal data (even with very few timepoints) can also be modeled with the intent of extracting a GRN (Lin et al., 2025).
The closest work to ours proves an identifiability bound for linear SDEs under a strong sparsity assumption (Dettling et al., 2023). However, their result doesn’t consider any interventional data, and the exact pattern of sparsity in the drift matrix is assumed to be known a priori, which is rarely the case in applied problems of interest. Another closely related work is Guan et al. (2024), which can simultaneously infer the drift and diffusion. However, this work applies only in an easier setting that assumes linearity, doesn’t study interventions, and assumes multiple temporal marginals rather than just the stationary distribution.
@@ 4 Main Results
We consider two main parameterizations of the drift as linear or a two-layer neural network (an MLP) to verify when the model may be uniquely identified up to appropriate invariances. Prior work has considered sparsity in the linear setting only (Dettling et al., 2023), but we focus on parameterizations that project to a low-dimensional space. In other words, although the ambient dimension $n$ may be large, we assume the hidden dimension $r$ is much smaller, at least $n>2r$, and ideally the number of interventions to uniquely identify the parameters should scale with $r$.
@@ 4.1 Linear case
We start with the linear case and consider the parameterization v(x)=(ABD)xv(x)=(AB-D)x where $A\in\mathbb{R}^{n\times r}$, $B\in\mathbb{R}^{r\times n}$, $D\in\mathbb{R}^{n\times n}$. Clearly $AB$ is a redundant parameterization of a rank $r$ matrix, but we use this notation to contrast the non-linear setting in Section 4.2. To ensure the drift is Hurwitz and the SDE has a stationary distribution, we assume $\|AB\|\leq\gamma<1$ and $D\succeq I$. Intuitively, $AB$ drive the dynamics while $D$ is a decay term to prevent unbounded dynamics. This parameterization is similar to the one used in Rohbeck et al. (2024), which also uses linear SDEs but considers hard interventions with having entirely new rows in the drift matrix, rather than shift interventions.
@@
In the linear setting, there exist choices for parameters $A,B,D$ and interventions $C$ such that the drift matrix ABDAB-D is not identifiable with less than nrn-r interventions.
The proof is provided in Appendix A. Intuitively, one can choose interventions that don’t affect the system dynamics at all. But this situation is pathological, and in the linear case under some weak distributional assumptions we can show identifiability.
@@
We take care to explain why these assumptions are not particularly restrictive. The low-rank constraint is already enforced by the drift function so $A$ and $B$ are almost surely full-rank when drawn from any density. The spectral norm bound is necessary to guarantee the SDE has a stationary distribution. And the rotational invariance assumption encodes an uninformative prior on which low-rank subspace governs the dynamics. Two simple choices that satisfy these assumptions are sampling $A$ and $B^{T}$ from either the uniform measure on the Stiefel manifold (rectangular matrices with orthonormal columns) or with iid Gaussian entries, and then scaling down to ensure a spectral norm strictly smaller than one, with $C$ sampled from any density.
Additionally, for theoretical tractability, we presume knowledge of the decay term $D$. This is a stronger assumption, but plausible in some applied contexts. For example, in single-cell genomics the decay rate of genes can be estimated with external experiments (e.g., from BRIC-seq (Imamachi et al., 2014)).
@@
The decay matrix $D$ is observed.
Altogether, we can now present a nearly tight identifiability result in the linear case.
@@
Consider the linear drift in Equation 5. Suppose $A$, $B$ are drawn from any density such that $\|A\|,\|B\|\leq\gamma<1$. Then the drift ABDAB-D is identifiable a.s. with $r$ interventions, and unidentifiable a.s. with at most r2r-2 interventions.
@@
If diagonal decay term $D$ is not inferred in advance, learning it simultaneously is comparable to the setting of robust PCA (Candès et al., 2011) or recovery of a diagonal plus a positive semidefinite low rank term (Saunderson et al., 2012). However, unlike the usual matrix completion setting where entries of the matrix are revealed uniformly at random, we have access to correlated low-rank measurements of the form $e_{i}c_{j}^{T}$ for $\{e_{i}\}_{i=1}^{n}$ the elementary basis. This setting is well studied with sub-Gaussian measurement vectors (Zhong et al., 2015), but the tools do not readily apply to our setting with non-random measurements and the constraints of the Lyapunov equation. Nevertheless, we observe in Section 5 that identifiability with an unknown diagonal decay matrix $D$ is empirically achieved, while still subject to the lower bound established in Theorem 4.4.
@@ 4.2 Nonlinear case
The nonlinear case is more challenging, because there is no longer a nearly closed form characterization of the stationary distribution. To apply tools from above, we consider when a linearized SDE can approximately capture the true dynamics, by restricting to contractive drift and small noise.
@@
The strongest constraint here is condition 1, upper bounding the first derivative of the activation, as it implies the noiseless dynamics are globally contractive, but it guarantees a stationary distribution exists for any intervention vector $c$. Furthermore, these conditions still allow for a wide class of possible activations $\sigma$, mainly constrained to be increasing with bounded first and second derivative. We observe the improved expressiveness of the stationary distribution of nonlinear SDEs in Figure 1. Note constraint 4 rules out functions that are linear (or locally linear). This enables a stronger possible identifiability guarantee than only recovering $A$ and $B$ up to rotation as in the linear case.
Figure 1: Contour plot of the stationary SDE under different activations and interventions. Activation contractivity enforces one mode, but the linear distribution is Gaussian with fixed covariance across interventions, while the nonlinear distribution can be more expressive.
@@
Setting $\epsilon=0$ in equation 5, the parameters are a.s. not identifiable with fewer than nrn-r interventions.
The proof is in Appendix A. Intuitively, the equilibrium points of the noiseless ODE approximate the mean of the SDE for small noise (note this is not true for larger noise, for example see Ma et al. (2015)). This suggests that in order to make use of the low-rank assumption on the dynamics, one must at least take advantage of second order moments of the SDE even in the small noise limit.
@@
Let $x^{*}$ be unique solution $v(x^{*})+c=0$, and define $L=Jv(x^{*})$. If $\omega$ solves the Lyapunov equation $L\omega+\omega L^{T}+I=0$, and $m_{\epsilon}$ and $\Sigma_{\epsilon}$ are the mean and covariance of the stationary distribution of Equation (5) intervened by $c$, we have for sufficiently small $\epsilon$,
@@
$\|\Sigma_{\epsilon}/\epsilon-\omega\|$ $\lesssim\frac{\epsilon^{1/2}r^{5/2}n^{1/2}M}{(1-\gamma)^{3}}.$ (7)
The proof is given in Appendix A. This is one instance of perturbation theory for SDEs (Gardiner, 2021; Sanz-Alonso & Stuart, 2016). Equipped with this fact, one can consider access to the stationary distribution of an SDE, specifically the first and second-order moments, and inspect what happens now as the noise goes to zero. In this case, we obtain a non-trivial identifiability bound:
@@
We give a brief sketch of the proof, to capture the novel elements and how the identifiability of $A$ and $B$ appears here but not in the linear case. The zero-noise limit approximately linearizes each intervened SDE. Unlike the exactly linear case, the stationary covariance is not identical across interventions. By Theorem 4.7, the covariance under the $i$th intervention approximately matches the SDE with drift $A(J\sigma(Bx_{i}^{*}))B$ where $x_{i}^{*}$ is the unique zero of the intervened drift $v(x)+c_{i}$.
Sufficient variability in these fixed points (guaranteed by our assumptions on $\sigma$) enables the identification of $A$ and $B$. Suppose one had access to the drift matrices: because the Jacobian term in the drift is diagonal, a linear combination of drifts across $r$ interventions can yield a rank-one matrix, which must correspond to the outer product of a column of $A$ and a row of $B$. But we don’t have access to the drift matrices directly, only the covariances they induce, and so the proof relies on finding a different set of matrices where rank-one matrices that characterize $A$ and $B$ can be isolated.
Crucially, the proof doesn’t require knowledge of the activation function itself. This enables identification of the bias term absorbed by $\sigma$, but furthermore one can identify the parameters $A$ and $B$ even if each elementwise activation in $\sigma$ is distinct. This observation motivates learnable activations, which we experiment with in Section 5.
We hypothesize the constraint that the SDE drift have a single absorbing point is unnecessary. If the ODE given by $\frac{dx}{dt}=v(x)+c_{i}$ had multiple stable equilibria, then in the limit of small noise, the SDE stationary distribution should approach a Gaussian mixture centered around each equilibrium point. The proof only requires certain linear independence conditions of vectors induced by each fixed point, so one intervention with multiple equilibria would offer essentially the same information as additional unimodal interventions. This is consistent with the observations in Lorch et al. (2024) that nonlinear SDE models generalize well with a limited number of interventions.